# MemoryBench: A Benchmark for Memory and Continual Learning in LLM Systems

### v2

@@ A.2.3 Probabilistic User Action Modeling
For the experiments in this paper, we instantiate these parameters by adopting statistical findings from large-scale user studies. Specifically, Shuster et al. (2022) report that about $6.5\%$ of users provide explicit feedback in the form of actions, and Chatterji et al. (2025) further show that such feedback exhibits an approximately $86\%$ to $14\%$ like-to-dislike ratio. Following these empirical findings, we set the overall user feedback rate to $6.5\%$ with the same like-to-dislike split. This yields target global probabilities of $P(L)\approx 0.0559$ and $P(D)\approx 0.0091$.
@@ A.3.1 Datasets Mixtures Details
For each dataset category, we randomly sampled up to 250 data points with a fixed random seed (if the dataset contained fewer than 250 samples, we used all available samples). The sampled data were split into training and test sets with a 4:1 ratio, and then merged across datasets to construct the final training and test sets for each category. Table 5 summarizes the datasets included in each domain, along with their total sizes and the number of sampled instances used in our experiments. Table 6 provides a similar overview for each task , showing how datasets are distributed and sampled within each category.
@@ A.3.2 Baselines
  • A-Mem (Xu et al., 2025a): An agentic memory system for LLM agents, designed to dynamically organize and evolve memories rather than relying on fixed structures. We adopt its official implementation with the provided APIs for adding and deleting memories, and use the default embedding model all-MiniLM-L6-v2. Following its implementation on the LoCoMo dataset, each message of a dialogue is stored as an independent memory entry, along with the dialogue ID and the message position.
  • Mem0 (Chhikara et al., 2025): A scalable memory-centric architecture that dynamically extracts, consolidates, and retrieves salient information from ongoing conversations. We use its open-source version with the provided APIs for memory addition and deletion, and employ Qwen3-Embedding-0.6B as the embedding model. According to the input forms of the API for memory addition, each full dialogue is added as a single memory entry. All memories, including both declarative memories from the dataset and procedural memories generated during interactions, are attributed to the same user identity.
  • MemoryOS (Kang et al., 2025): A memory operating system that introduces a hierarchical and comprehensive memory management framework for LLM agents, inspired by operating system principles. We adopt its official implementation with the provided API for adding memories and use all-MiniLM-L6-v2 as the embedding model. As required by its API, we combine one user message with its corresponding assistant reply into a single memory entry. All memories are attributed to the same user identity. In particular, for compatibility with Chinese datasets, we translated several internal English prompts into Chinese when used in Chinese datasets.
@@ A.3.4 Off-Policy Experiments
We constructed multiple LLMsys using Qwen3-8B as the backbone model and evaluated their performance on MemoryBench. In the off-policy experiments, the LLMsys are required to leverage the dialogues generated on the training set in order to better solve problems in the test set. For each domain and task, we aggregate the results by reporting both the min-max normalization score and the Z-score. The results of min-max normalization scores and z-score scores are shown in Table 7 and Table 8 respectively. To enable a cross-experimental comparison, the min and max values used for min-max normalization in all other experiments are aligned with those obtained from the Qwen3-8B off-policy experiments on each dataset.
@@
Vanilla 0.5259 0.7152 0.4781 0.4168 0.7792 0.6258 0.7442
BM25-M 0.5990 0.6994 0.4913 0.3946 0.7537 0.6325 0.7403
BM25-S 0.6302 0.7195 0.4754 0.4093 0.7757 0.6280 0.736
Embed-M 0.6194 0.6995 0.4521 0.4427 0.7301 0.5969 0.7226
Embed-S 0.6537 0.7066 0.4759 0.4545 0.7653 0.6316 0.7516
A-Mem 0.6170 0.6908 0.4698 0.4020 0.7522 0.6059 0.7278
Mem0 - 0.6617 0.4279 - 0.7353 0.5869 0.6496
MemoryOS 0.6011 0.6428 0.4336 0.3449 0.7780 0.4974 0.7241
@@
Vanilla -0.5096 0.1373 0.0528 0.0410 0.1305 0.1117 0.0959
BM25-M -0.0700 0.0380 0.1166 -0.0434 -0.0292 0.1136 -0.0209
BM25-S 0.1013 0.1467 0.0353 0.0269 0.1066 0.1167 0.0833
Embed-M -0.0600 0.0147 -0.0397 0.0938 -0.1782 -0.0068 -0.0112
Embed-S 0.0951 0.0881 0.0554 0.1363 0.0446 0.1353 0.1494
A-Mem -0.0134 -0.0576 0.0505 -0.0141 -0.0386 0.0213 0.0108
Mem0 - -0.1840 -0.1491 - -0.1566 -0.0498 -0.3244
MemoryOS -0.0531 -0.1832 -0.1217 -0.1994 0.1210 -0.4420 0.0172
To provide a complete view, we also present the detailed evaluation results before aggregation in Table 9.
@@
Open LoCoMo F1 0.3717 0.4252 0.2589 0.5349 0.5184 0.4355 - 0.4515
DialSim-friends Accuracy 0.1765 0.1176 0.2353 0.3529 0.2941 0.2941 - 0.1176
DialSim-bigbang Accuracy 0.3529 0.1765 0.1765 0.1176 0.3529 0.0588 - 0.0588
DialSim-theoffice Accuracy 0.3529 0.0588 0.4118 0.2353 0.2941 0.2353 - 0.0588
HelloBench-C.D. Avg. Score 0.8071 0.7239 0.7832 0.7381 0.7821 0.7846 - 0.7846
WritingPrompts Meteor 0.2345 0.2160 0.2396 0.2051 0.2274 0.2162 - 0.2241
WritingBench-C.D. Score 6.7326 6.3372 6.8256 6.4186 6.4884 6.4884 - 6.2093
NFCats Score 4.6800 4.7200 4.6000 4.3800 4.5000 4.4800 - 4.6000
Academic
HelloBench-
A.K.-QA
Avg. Score 0.8340 0.8776 0.8636 0.8239 0.8214 0.8526 0.8063 0.8271
@@
Avg. Score 0.8250 0.8495 0.8289 0.8328 0.8441 0.8422 0.8167 0.7637
IdeaBench BERTScore 0.5632 0.5549 0.5676 0.555 0.5701 0.5536 0.5435 0.5608
@@
LimitGen-Syn Accuracy 0.5200 0.5800 0.5000 0.5800 0.5400 0.5000 0.6000 0.2400
LimitGen-Syn Rating 1.3800 1.5800 1.5200 1.6600 1.5400 1.4600 1.3200 0.5400
WritingBench-A.E. Score 6.9118 6.4412 7.0000 6.5294 6.5294 7.0294 6.6471 5.5000
@@
0.7199 0.7134 0.7037 0.7146 0.7253 0.7377 0.7072 0.7439
Penalcode
Index F1
0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000
Time Score 0.7289 0.6843 0.6570 0.7161 0.7082 0.7190 0.7023 0.7112
Amount Score 0.4770 0.4113 0.4557 0.4713 0.4342 0.4610 0.4511 0.4374
@@
ROUGE-L 0.0875 0.0943 0.0862 0.0463 0.0778 0.0468 0.0216 0.0441
LexEval-QA ROUGE-L 0.1079 0.1453 0.1184 0.1151 0.1192 0.1320 0.1121 0.1261
WritingBench-P.L. Score 7.0732 6.7561 7.0000 6.5610 6.6585 6.9512 6.9268 5.5610
@@
Vanilla 0.5708 -0.3272 0.7200 0.1427 0.4771 0.0732
BM25-M 0.5250 -0.4915 0.7221 0.1756 0.5204 0.2568
BM25-S 0.5571 -0.3834 0.7343 0.2480 0.4996 0.1744
Embed-M 0.5240 -0.5263 0.7192 0.1412 0.4792 0.0939
Embed-S 0.5800 -0.3367 0.7327 0.2456 0.4825 0.0980
A-Mem 0.5114 -0.5318 0.7388 0.2448 0.4822 0.1068
Mem0 - - 0.6993 -0.0021 0.4782 0.0816
MemoryOS 0.4961 -0.6334 0.6484 -0.1877 0.4259 -0.1588

### v4

@@ A.2.3 Probabilistic User Action Modeling
For the experiments in this paper, we instantiate these parameters by adopting statistical findings from large-scale user studies. Specifically, shuster2022blenderbot report that about $6.5\%$ of users provide explicit feedback in the form of actions, and chatterji2025people further show that such feedback exhibits an approximately $86\%$ to $14\%$ like-to-dislike ratio. Following these empirical findings, we set the overall user feedback rate to $6.5\%$ with the same like-to-dislike split. This yields target global probabilities of $P(L)\approx 0.0559$ and $P(D)\approx 0.0091$.
@@ A.2.4 Prompt Templates
A.2.5 Human Annotation of Feedback Generation Quality
To better evaluate the quality of simulated user feedback in MemoryBench, we randomly sampled 200 tasks across 4 task format categories, 3 domains, and 2 languages together with feedback generated by the LLM-as-User-Simulator from the training set in the off-policy experimental setting for human annotation. The distribution of the sampled task cases is shown in Table 6. For each task case, we conducted a two-step human annotation. First, we hired professional human annotators to write verbose feedback for each task based on the task needs and evaluation criteria. Second, we conducted crowdsourcing by recruiting 9 annotators (mostly undergraduate and graduate students on campus) to conduct pairwise annotations to compare the LLM-simulated user feedback with human feedback in terms of language naturalness, relevance to the task context, and which looks more like a feedback written by human users:
  • Naturalness: Whether the feedback sounds like something a real user would naturally say, using authentic vocabulary, tone, and flow without unnatural repetition or stiffness.
  • Relevance: Whether the feedback stays on-topic, be logically coherent, and provide specific, valuable guidance or information related to the conversation.
  • Overall Quality: Considering both naturalness and relevance, whether the feedback looks like a plausible, reasonable response from a real user in the given context.
Each task case is annotated by three different annotators. The Fleiss’ Kappa of annotators as well as the annotation results are shown in Table 7. As shown in the table, the Kappa value of Naturalness is low, which means that it was difficult for the annotators to consistently distinguish whether the human-written or LLM-simulated user feedback is more natural in language. This is not surprising as modern LLMs have already shown strong abilities in high-quality text generation. The Kappa value of Relevance and Overall Quality is better, but still not high. This indicates that, in general, it is difficult to distinguish human-written verbose feedback with simulated verbose feedback generated by the LLM-as-user framework in MemoryBench. This is further supported by the final annotation results. In Table 7, half of the annotators believe that the simulated feedback is more natural than human-written feedback, and a slightly more than half of the annotators feel that the simulated feedback is better than the human-written feedback in terms of relevance and overall quality. These are supportive evidences showing that our LLM-as-user framework can generate high-quality user feedback that are indistinguishable from real human feedback.
Table 6: Distribution of sampled task cases for human annotation.
Divided by Language Task Domain
English Chinese LiSo SiLo LiLo SiSo Open Academic Legal
Samples 131 69 43 46 74 37 66 77 57
Table 7: Human annotation results comparing the simulated and human-written feedback.
Annotator Kappa Simulated Feedback Win-rate Human Feedback Win-rate
Naturalness 0.093 50.0% 50.0%
Relevance 0.380 59.5% 40.5%
Overall Quality 0.273 59.0% 41.0%
@@ A.3.1 Datasets Mixtures Details
For each dataset category, we randomly sampled up to 250 data points with a fixed random seed (if the dataset contained fewer than 250 samples, we used all available samples). The sampled data were split into training and test sets with a 4:1 ratio, and then merged across datasets to construct the final training and test sets for each category. Table 8 summarizes the datasets included in each domain, along with their total sizes and the number of sampled instances used in our experiments. Table 9 provides a similar overview for each task , showing how datasets are distributed and sampled within each category. We also provide an extended version of MemoryBench using all data from all datasets with the same training and test sets sampling ratio as MemoryBench-Full [^5] https://huggingface.co/datasets/THUIR/MemoryBench-Full. However, the sizes of different datasets in MemoryBench-Full vary significantly, which may require special handling when calculating averages to prevent any single dataset from disproportionately influencing the results, so we do not include the experiments on it in this paper.
@@ A.3.2 Baselines
  • A-Mem (xu2025mem): An agentic memory system for LLM agents, designed to dynamically organize and evolve memories rather than relying on fixed structures. A-Mem consists of three major modules: memory note construction, dynamic memory linking for continual evolution, and memory retrieval. Specifically, A-Mem uses an LLM to extract a structured memory note for each input text and encodes these notes into embeddings. It then uses embedding similarity to retrieve the most relevant existing memories, constructs memory links between them, and updates linked memories jointly via LLM prompting. During inference, the most relevant memory notes are retrieved and subsequently incorporated into the context to augment the model’s generation. We adopt its official implementation with the provided APIs for adding and deleting memories, and use the default embedding model all-MiniLM-L6-v2. Following its implementation on the LoCoMo dataset, each message of a dialogue is stored as an independent memory entry, along with the dialogue ID and the message position.
  • Mem0 (chhikara2025mem0): A scalable memory-centric architecture that dynamically extracts, consolidates, and retrieves salient information from ongoing conversations. Mem0 includes two core modules: memory extraction and memory update. For extraction, Mem0 uses dialogue summaries and recent messages to generate candidate memories. For update, Mem0 leverages the LLM’s inherent reasoning capability for efficient memory maintenance. The LLM compares new information against existing memories and chooses among four explicit operations—ADD, UPDATE, DELETE, or NOOP—to guarantee data consistency and minimize redundancy in the memory store. Mem0 also proposes an extension, Mem0g, which structures memories as a knowledge graph; however, Mem0g is not used in our experiments. We use its open-source version with the provided APIs for memory addition and deletion, and employ Qwen3-Embedding-0.6B as the embedding model. According to the input forms of the API for memory addition, each full dialogue is added as a single memory entry. All memories, including both declarative memories from the dataset and procedural memories generated during interactions, are attributed to the same user identity.
  • MemoryOS (kang2025memory): A memory operating system that introduces a hierarchical and comprehensive memory management framework for LLM agents, inspired by operating system principles. MemoryOS includes four core modules: memory storage, update, retrieval, and generation. Analogous to an operating system, it organizes memories into short-term, mid-term, and long-term storage units, and refreshes them through a segmented paging architecture driven by dialogue-chain dynamics and heat-based mechanisms. During retrieval, MemoryOS performs semantic segmentation to query memories across different layers, and integrates the retrieved segments into a coherent prompt that supports downstream generation. We adopt its official implementation with the provided API for adding memories and use all-MiniLM-L6-v2 as the embedding model. As required by its API, we combine one user message with its corresponding assistant reply into a single memory entry. All memories are attributed to the same user identity. In particular, for compatibility with Chinese datasets, we translated several internal English prompts into Chinese when used in Chinese datasets.
@@ A.3.4 Off-Policy Experiments
We constructed multiple LLMsys using Qwen3-8B as the backbone model and evaluated their performance on MemoryBench. To generate the training dialogues used in the off-policy setting, we employed Qwen3-32B as the feedback simulator model and Qwen3-8B as the assistant model. In the off-policy experiments, the LLMsys are required to leverage the dialogues generated on the training set in order to better solve problems in the test set. For each domain and task, we aggregate the results by reporting both the min-max normalization score and the Z-score. The results of min-max normalization scores and z-score scores are shown in Table 10 and Table 11 respectively. To enable a cross-experimental comparison, the min and max values used for min-max normalization in all other experiments are aligned with those obtained from the Qwen3-8B off-policy experiments on each dataset.
@@
Vanilla 0.6523 0.7154 0.4637 0.4091 0.7517 0.6417 0.7302
BM25-M 0.5720 0.6931 0.5011 0.3869 0.7287 0.6318 0.7275
BM25-S 0.6503 0.7110 0.4776 0.4560 0.7260 0.6310 0.7330
Embed-M 0.5760 0.7051 0.4635 0.3844 0.7120 0.5908 0.6919
Embed-S 0.6582 0.7185 0.4806 0.4661 0.7325 0.6328 0.7305
A-Mem 0.6243 0.7106 0.4684 0.4303 0.7172 0.6139 0.7147
Mem0 - 0.6965 0.4451 - 0.7305 0.5957 0.6745
MemoryOS 0.5894 0.6212 0.4290 0.3711 0.7493 0.4879 0.7452
@@
Vanilla 0.1583 0.1450 -0.0019 0.0162 -0.0467 0.1828 0.0392
BM25-M -0.1813 0.0293 0.1531 -0.0375 -0.1930 0.1166 -0.0641
BM25-S 0.1030 0.1265 0.0496 0.1201 -0.2169 0.1339 0.0855
Embed-M -0.1579 0.0760 0.0081 -0.0514 -0.2970 -0.0343 -0.1608
Embed-S 0.1081 0.1775 0.0669 0.1629 -0.1718 0.1373 0.0495
A-Mem 0.0340 0.1124 0.0458 0.0743 -0.2650 0.0555 -0.0814
Mem0 - 0.0045 -0.0843 - -0.1836 -0.0049 -0.2593
MemoryOS -0.0875 -0.2621 -0.1561 -0.0824 -0.0619 -0.4876 0.1074
To provide a complete view, we also present the detailed evaluation results before aggregation in Table 12 and the variance in Table 13&14.
@@
Open LoCoMo F1 0.3759 0.3795 0.4280 0.3056 0.5730 0.4148 - 0.4595
DialSim-friends Accuracy 0.1765 0.1765 0.4118 0.2941 0.1765 0.2941 - 0.0588
DialSim-bigbang Accuracy 0.3529 0.1176 0.2941 0.0588 0.3529 0.0588 - 0.0588
DialSim-theoffice Accuracy 0.3529 0.1176 0.3529 0.2353 0.4118 0.2353 - 0.0000
HelloBench-C.D. Avg. Score 0.8541 0.7220 0.8098 0.7892 0.8135 0.8223 - 0.7910
WritingPrompts Meteor 0.2337 0.2219 0.2356 0.2037 0.2284 0.2181 - 0.2270
WritingBench-C.D. Score 6.6860 6.4419 6.5233 6.4651 6.6279 6.6047 - 6.3372
NFCats Score 4.5800 3.9800 4.3600 4.1200 4.2400 4.5000 - 4.3200
Academic HelloBench-A.K.-QA Avg. Score 0.8364 0.8599 0.8423 0.8243 0.8724 0.8653 0.8451 0.8592
@@
Avg. Score 0.8475 0.8353 0.8544 0.8544 0.8795 0.8446 0.8225 0.8392
IdeaBench BERTScore 0.5583 0.5590 0.5654 0.5598 0.5656 0.5568 0.5451 0.5585
@@
LimitGen-Syn Accuracy 0.4600 0.4800 0.4000 0.5200 0.4800 0.4600 0.5600 0.1200
Rating 1.3800 1.2400 1.3200 1.5400 1.5200 1.3800 1.5200 0.3200
WritingBench-A.E. Score 6.9118 6.5588 6.8529 6.5000 6.5000 6.8824 6.8235 4.9118
@@
0.7293 0.7139 0.6964 0.7287 0.7213 0.7444 0.7471 0.7016
Time Score 0.7312 0.7009 0.6554 0.7005 0.6939 0.7060 0.6874 0.7026
Amount Score 0.4506 0.4224 0.4247 0.4437 0.4183 0.4892 0.4815 0.4796
@@
ROUGE-L 0.0644 0.1116 0.0864 0.0561 0.0863 0.0461 0.0461 0.0470
LexEval-QA ROUGE-L 0.0986 0.1359 0.1284 0.1240 0.1175 0.1335 0.1147 0.1009
WritingBench-P.L. Score 7.0732 6.8293 6.7805 6.7561 6.9268 6.7561 7.0000 5.9268
Table 13: Detailed evaluation results of LLMsys with Qwen3-8B backbone in the off-policy setting with both raw performance scores using their dataset-specific evaluation metrics where the value following the ± symbol represents the standard deviation across the test data. Except for metrics marked with $\downarrow$, higher values indicate better performance.
Domain Dataset Metric Vanilla A-Mem Mem0 MemoryOS
Open LoCoMo F1 0.3759±0.4232 0.4148±0.4709 - 0.4595±0.4846
DialSim-friends Accuracy 0.1765±0.3812 0.2941±0.4556 - 0.0588±0.2353
DialSim-bigbang Accuracy 0.3529±0.4779 0.0588±0.2353 - 0.0588±0.2353
DialSim-theoffice Accuracy 0.3529±0.4779 0.2353±0.4242 - 0.0000±0.0000
HelloBench-C.D. Avg. Score 0.8541±0.1301 0.8223±0.1490 - 0.7910±0.2006
WritingPrompts Meteor 0.2337±0.0434 0.2181±0.0588 - 0.2270±0.0503
WritingBench-C.D. Score 6.6860±1.2691 6.6047±1.4409 - 6.3372±1.5296
NFCats Score 4.5800±0.8964 4.5000±0.8307 - 4.3200±0.8352
Academic HelloBench-A.K.-QA Avg. Score 0.8364±0.1162 0.8653±0.0837 0.8451±0.1265 0.8592±0.1282
HelloBench-
A.K.-Writing
Avg. Score 0.8475±0.0654 0.8446±0.0738 0.8225±0.1062 0.8392±0.0775
IdeaBench BERTScore 0.5583±0.0336 0.5568±0.0338 0.5451±0.0330 0.5585±0.0340
LLM Rating
Score
4.3000±2.1932 3.8000±2.1541 3.6600±1.9658 4.4400±2.1369
LLM Novelty
Ranking Score
0.6600±0.4082 0.5933±0.4232 0.5333±0.4570 0.4933±0.4333
LLM Feasibility
Ranking Score
0.1867±0.3138 0.1467±0.2918 0.0600±0.2180 0.1200±0.2473
SciTechNews ROUGE-L 0.1222±0.0213 0.1149±0.0278 0.1137±0.0194 0.0900±0.0273
BERTScore-F1 0.8179±0.0090 0.8173±0.0119 0.8183±0.0105 0.8110±0.0121
[…]
LimitGen-Syn Accuracy 0.4600±0.4984 0.4600±0.4984 0.5600±0.4964 0.1200±0.3250
Rating 1.3800±1.5085 1.3800±1.5861 1.5200±1.4730 0.3200±0.9042
WritingBench-A.E. Score 6.9118±1.8209 6.8824±1.6227 6.8235±1.6174 4.9118±2.8735
Legal JuDGE
Reasoning
Meteor
0.5138±0.1658 0.4604±0.1511 0.4823±0.1477 0.4770±0.1785
Judge
Meteor
0.4077±0.2225 0.4364±0.2364 0.3559±0.1600 0.3150±0.1528
Reasoning
BERTScore
0.8155±0.0529 0.7601±0.0876 0.8073±0.0452 0.8092±0.0639
Judge
BERTScore
0.7897±0.0772 0.7971±0.0831 0.7693±0.0629 0.7308±0.1185
Crime
Recall
0.9800±0.1400 1.0000±0.0000 0.9600±0.1960 1.0000±0.0000
Crime
Precision
0.9600±0.1685 0.9700±0.1187 0.9500±0.2062 0.9700±0.1187
Penalcode
Index Recall
0.7615±0.2329 0.7352±0.2522 0.7130±0.2633 0.7195±0.2741
Penalcode
Index Precision
0.7293±0.2485 0.7444±0.2608 0.7471±0.2592 0.7016±0.2844
Time Score 0.7312±0.2297 0.7060±0.2681 0.6874±0.2444 0.7026±0.2446
Amount Score 0.4506±0.3583 0.4892±0.3642 0.4815±0.3583 0.4796±0.3783
LexEval-Summarization ROUGE-L 0.2206±0.0894 0.2299±0.0882 0.1961±0.0785 0.2212±0.0978
LexEval-Judge ROUGE-L 0.0644±0.1261 0.0461±0.1068 0.0461±0.0946 0.0470±0.1104
LexEval-QA ROUGE-L 0.0986±0.0710 0.1335±0.0614 0.1147±0.0625 0.1009±0.0691
WritingBench-P.L. Score 7.0732±1.3505 6.7561±1.3576 7.0000±1.4314 5.9268±2.7265
Table 14: The continued part of Table 13
Domain Dataset Metric BM25-M BM25-S Embed-M Embed-S
Open LoCoMo F1 0.3795±0.4557 0.4280±0.4311 0.3056±0.3816 0.5730±0.4358
DialSim-friends Accuracy 0.1765±0.3812 0.4118±0.4922 0.2941±0.4556 0.1765±0.3812
DialSim-bigbang Accuracy 0.1176±0.3222 0.2941±0.4556 0.0588±0.2353 0.3529±0.4779
DialSim-theoffice Accuracy 0.1176±0.3222 0.3529±0.4779 0.2353±0.4242 0.4118±0.4922
HelloBench-C.D. Avg. Score 0.7220±0.2525 0.8098±0.1787 0.7892±0.2130 0.8135±0.1555
WritingPrompts Meteor 0.2219±0.0525 0.2356±0.0463 0.2037±0.0582 0.2284±0.0481
WritingBench-C.D. Score 6.4419±1.8018 6.5233±1.3786 6.4651±1.8783 6.6279±1.5704
NFCats Score 3.9800±1.2883 4.3600±1.0151 4.1200±1.0889 4.2400±1.0688
Academic HelloBench-A.K.-QA Avg. Score 0.8599±0.0870 0.8423±0.0935 0.8243±0.1835 0.8724±0.0781
HelloBench-
A.K.-Writing
Avg. Score 0.8353±0.0633 0.8544±0.0618 0.8544±0.0681 0.8795±0.0594
IdeaBench BERTScore 0.5590±0.0392 0.5654±0.0367 0.5598±0.0376 0.5656±0.0357
LLM Rating
Score
4.3400±2.0553 4.3000±2.2023 4.2000±2.1633 4.3061±2.2150
LLM Novelty
Ranking Score
0.5667±0.3844 0.6400±0.4155 0.5867±0.4246 0.5667±0.4435
LLM Feasibility
Ranking Score
0.1200±0.2473 0.2000±0.3333 0.1600±0.2924 0.1333±0.3266
SciTechNews ROUGE-L 0.1210±0.0308 0.1209±0.0231 0.1236±0.0283 0.1242±0.0251
BERTScore-F1 0.8185±0.0105 0.8187±0.0104 0.8181±0.0139 0.8155±0.0104
[…]
LimitGen-Syn Accuracy 0.4800±0.4996 0.4000±0.4899 0.5200±0.4996 0.4800±0.4996
Rating 1.2400±1.4221 1.3200±1.7713 1.5400±1.5901 1.5200±1.6400
WritingBench-A.E. Score 6.5588±2.0174 6.8529±1.5743 6.5000±1.8511 6.5000±2.0037
Legal JuDGE
Reasoning
Meteor
0.4943±0.1585 0.4996±0.1657 0.4918±0.1688 0.5025±0.1539
Judge
Meteor
0.4354±0.2211 0.3425±0.1785 0.4082±0.2172 0.3868±0.2012
Reasoning
BERTScore
0.8134±0.0504 0.8061±0.0627 0.8015±0.0615 0.8163±0.0428
Judge
BERTScore
0.7822±0.1361 0.7410±0.1264 0.7692±0.1359 0.7710±0.0701
Crime
Recall
0.9600±0.1960 0.9800±0.1400 0.9800±0.1400 0.9800±0.1400
Crime
Precision
0.9500±0.2062 0.9600±0.1685 0.9600±0.1685 0.9600±0.1685
Penalcode
Index Recall
0.7504±0.2618 0.7836±0.2425 0.7402±0.2535 0.7340±0.2257
Penalcode
Index Precision
0.7139±0.2466 0.6964±0.2466 0.7287±0.2324 0.7213±0.2522
Time Score 0.7009±0.2620 0.6554±0.2599 0.7005±0.2603 0.6939±0.2354
Amount Score 0.4224±0.3542 0.4247±0.3827 0.4437±0.3712 0.4183±0.3532
LexEval-Summarization ROUGE-L 0.2312±0.0944 0.2259±0.0941 0.2228±0.1000 0.2271±0.0830
LexEval-Judge ROUGE-L 0.1116±0.1535 0.0864±0.1488 0.0561±0.1158 0.0863±0.1473
LexEval-QA ROUGE-L 0.1359±0.0594 0.1284±0.0719 0.1240±0.0812 0.1175±0.0634
WritingBench-P.L. Score 6.8293±1.6064 6.7805±1.5852 6.7561±1.9099 6.9268±1.1974
@@
Vanilla 0.6778 0.2917 0.7297 0.2205 0.4630 -0.0061
BM25-M 0.6521 0.1743 0.7264 0.2352 0.4950 0.1504
BM25-S 0.7045 0.3289 0.7384 0.2593 0.4874 0.1250
Embed-M 0.6325 0.0809 0.7365 0.2366 0.4663 0.0342
Embed-S 0.6964 0.3149 0.7369 0.2736 0.4915 0.1265
A-Mem 0.6302 0.1041 0.7197 0.1557 0.4943 0.1497
Mem0 - - 0.6896 -0.0307 0.4739 0.0465
MemoryOS 0.6047 -0.0293 0.6450 -0.2241 0.4541 -0.0247